# Multicriteria Optimization and Decision Making

### v1

@@
Multicriteria Optimization and Decision Making
Principles, Algorithms and Case Studies
(Michael Emmerich and André Deutz
Lecture Notes
MSc Course 2012-2023, LIACS, Leiden University
The Netherlands
Updated and Revised Edition, June 2024 [Uncaptioned image] )
@@
For several reasons multicriteria optimization and decision making is an exciting field of computer science and operations research. Part of its fascination stems from the fact that in MCO and MCDM different scientific fields are addressed. Firstly, to develop the general foundations and methods of the field one has to deal with structural sciences, such as algorithmics, relational logic, operations research, and numerical analysis:
@@
  • What are the essential differences between single objective and multiobjective optimization?
  • How can we rank solutions? What different types of orderings are used in decision theory and how are they related to each other?
  • Given a decision model or optimization problem, which formal conditions need to be satisfied for solutions to be optimal?
  • How can we construct algorithms that obtain optimal solutions, or approximations to them, in an efficient way?
@@
Whenever it comes to decision making in the real world, these decisions will be made by people responsible for it. In order to understand how people come to decisions and how the psychology of individuals (cognition, individual decision making) and organizations (group decision making) needs to be studied. Questions like the following may arise:
@@
  • Which different strategies are used by people to come to decisions? How can satisfaction be measured? What strategies are promising in obtaining satisfactory decisions?
@@
  • How do groups of people come to decisions? What are conflicts and how can they be avoided? How to deal with minority interests in a democratic decision process? Can these aspects be integrated into formal decision models?
Moreover, decisions are always related to some real world problem. Given an application field, we may find very specific answers to the following questions:
@@
  • By which means can we retrieve the values for the criteria (experiments, surveys, function evaluations)? Are there any particular problems with these measurements (dangers, costs), and how to deal with them? What are the uncertainties in these measurements?
@@
  • What are typical decision processes in the field, and what implications do they have for the design of decision support systems?
@@ 1.1 Viewing mulicriteria optimization as a task in system design and analysis
The discussion above can be seen as a rough sketch of questions that define the scope of multicriteria optimization and decision making. However, it needs to be clarified more precisely what is going to be the focus of these lecture notes. For this reason we want to approach the problem class from the point of view of system design and analysis. Here, with system analysis, we denote the interdisciplinary research field, that deals with the modeling, simulation, and synthesis of complex systems.
Beside experimentation with a physical system, often a system model is used. Nowadays, system models are typically implemented as computer programs that solve (differential) equation systems, simulate interacting automata, or stochastic models. We will also refer to them as simulation models. An example for a simulation model based on differential equations would be the simulation of the fluid flow around an airfoil based on the Navier Stokes equations. An example for a stochastic system model, could be the simulation of a system of elevators, based on some agent based stochastic model.
@@
In Figure 1.1 different tasks of systems analysis based on simulation models are displayed in a schematic way. Modeling means to identify the internal structure of the simulation model. This is done by looking at the relationship between known inputs and outputs of the system. In many cases, the internal structure of the system is already known up to a certain granularity and only some parameters need to be identified. In this case we usually speak of calibration of the simulation model instead of modeling. In control theory, also the term identification is common.
Once a simulation-model of a system is given, we can simulate the system, i.e. predict the state of the output variables for different input vectors. Simulation can be used for predicting the output for not yet measured input vectors. Usually such model-based predictions are much cheaper than to do the experiment in the real world. Consider for example crash test simulations or the simulation of wind channels. In many cases, such as for future predictions, where time is the input variable, it is even impossible to do the experiments in the physical world. Often the purpose of simulation is also to learn more about the behavior of the systems. In this case systematic experimenting is often used to study effects of different input variables and combinations of them. The field of Design and Analysis of Computer Experiments (DACE) is devoted to such systematic explorations of a systems behavior.
Finally, we may want to optimize a system: In that case we basically specify what the output of the system should be. We also are given a simulation-model to do experiments with, or even the physical system itself. The relevant, open question is how to choose the input variables in order to achieve the desired output. In optimization we typically want to maximize (or minimize) the value of an output variable.
@@ 1.2 Formal Problem Definitions in Mathematical Programming
The objective function $f$ is a function to be minimized (or maximized [^1] Maximization can be rewritten as minimization by changing the sign of the objective function, that is, replacing $f(\mathbf{x})\rightarrow\max$ with $-f(\mathbf{x})\rightarrow\min$). This is the goal of the optimization. The function $f$ can be evaluated for every point $\mathbf{x}$ in the search space (or decision space). Here the search space is defined by a set of intervals, that restrict the range of variables, so called bounds or box constraints. Besides this, variables can be integer variables that is they are chosen from $\mathbb{Z}$ or subsets of it, or continuous variable (from $\mathbb{R}$). An important special case of integer variables are binary variables which are often used to model binary decisions in mathematical programming.
Whenever inequality and equality constraints are stated explicitly, the search space $\mathcal{X}$ can be partitioned in a feasible search space $\mathcal{X}_{f}\subseteq\mathcal{X}$ and an infeasible subspace $\mathcal{X}-\mathcal{X}_{f}$. In the feasible subspace all conditions stated in the mathematical programming problem are satisfied. The conditions in the mathematical program are used to avoid constraint violations in the system under design, e.g., the excess of a critical temperature or pressure in a chemical reactor (an example for an inequality constraint), or the keeping of conservation of mass formulated as an equation (an example for an equality constraint). The conditions are called constraints. Due to a convention in the field of operations research, constraints are typically written in a standardized form such that 00 appears on the right hand side side. Equations can easily be transformed into the standard form by means of algebraic equivalence transformations.
Based on this very general problem definition we can define several classes of optimization problems by looking at the characteristics of the functions $f$, $g_{i},i=1,\dots,n_{g}$, and $h_{i},i=1,\dots,n_{h}$. Some important classes are listed in Table 1.1.
@@ 1.2.1 Other Problem Classes in Optimization
There are also other types of mathematical programming problems. These are, for instance based on:
@@
In some cases, the demand that a solution consists of a vector of variables is too restrictive and instead we can define the search space as some set $\mathcal{X}$. In order to capture also these kind of problems a more general definition of a general optimization problem can be used:
@@
Another important special case is given, if $\mathcal{X}=\mathbb{R}^{n}$. Such problems are defined as continuous unconstrained optimization problems or, simply, unconstrained optimization problems.
For notational convenience in the following we will refer mostly to the generic definition of an optimization problem given in equation 1.10, whenever constraint treatment is not particularly addressed. In such cases we assume that $\mathcal{X}$ already contains only feasible solutions.
@@ 1.2.2 Multiobjective Optimization
Before proceeding in this direction, it is however important to note, that many difficulties of solving single objective optimization problems are inherited by the more general class of multiobjective optimization problems. We will therefore first summarize these.
@@ 1.3.1 Problem Difficulty in Continuous Optimization
In continuous optimization the metaphor of a optimization landscape is often used in order to define problem difficulty. As opposed to just talking about a function, when using the term (search) landscapes one explicitly requires the search space to be equipped with a neighborhood structure, which could be a metric or a topology. This topology is typically the standard topology on $\mathbb{R}^{n}$ and as a metric typically the Euclidean metric is used.
As we shall discuss with more rigor in Chapter 4, this gives rise to definitions such as local optima, which are points that cannot be improved by replacing them with neighboring points. For many optimum-seeking algorithms, it is difficult to escape from such points or find a good direction (in case of plateaus). If local optima are not also global optima the algorithm might return a suboptimal solutions.
@@
Finally, discontinuities and ruggedness of the landscape make problems difficult to solve for many solvers. Discontinuities are abrupt changes of the function value in some neighborhood. In particular these cause difficulties for optimization methods that assume the objective function to be continuous, that is they assume that similar inputs cause similar outputs. A common definition of continuity is that of Lipschitz continuity.
@@
Besides continuity assumptions, also differentiability of the objective function and constraints, convexity and mild forms of non-linearity (as given in convex quadratic optimization), as well as limited interaction between variables can make a continuous problem easier to solve. The degree of interaction between variables is given by the number of variables in the term of the objective function: Assume it is possible to (re)-write the optimization problem in the form $\sum_{i=1}^{n}f_{i}(x_{i_{1}},\dots,x_{i_{k(i)}})\rightarrow\max$, then the value of $k(i)$ is the degree of interaction in the $i$-th component of the objective function. In case of continuous objective functions it can be shown that problems with a low degree of interaction can be solved more efficiently in terms of worst case time complexity[69]. One of the reasons why convex quadratic problems can be solved efficiently is that, given the Hessian matrix, the coordinate system can be transformed by simple rotation in such a way that these problems become decomposable, i.e. $k(i)$ is bounded by $1$.
@@ 1.3.2 Problem Difficulty in Combinatorial Optimization
Having said this, a combinatorial optimization problem can be defined by means of a pseudo-boolean objective function, i.e. $f:\{0,1\}^{n}\rightarrow\mathbb{R}$ and stating the goal $f(\mathbf{x})\rightarrow\min$. Theoretical computer science has developed a rich theory on the complexity of decision problems. A decision problem is the problem of answering a query on input of size $n$ with the answer being either yes or no. In order to relate the difficulty of optimization problems to the difficulty of decision problems it is beneficial to formulate so-called decision versions of optimization problems.
@@
A decision problem is said to belong to the class P if there exists an algorithm on a Turing machine222or any in any common programming language operating on infinite memory and not using parallel processing and not assuming constant time complexity for certain infinite precision floating point operations such as the floor function. that solves it with a time complexity that grows at most polynomially with the size $n$ of the input. It belongs to the class NP if a candidate solution $\mathbf{x}$ of size $n$ can be verified (’checked’) with polynomial time complexity (does it satify the formula $f(\mathbf{x})\leq k$ or not). Obviously, the class NP subsumes the class P, but the P perhaps not neccessarily subsumes NP. In fact, the question whether P subsumes NP is the open problem often discussed in theoretical computer science known as the problem ’P$=$NP’. Under the assumption ’P $\neq$ NP’, that is that P does include NP, it is meaningful to introduce the complexity class of NP complete problems:
@@
A decision problem $\ textrm{D}$ is NP-complete if it belongs to NP, and every instance of another problem in NP can be transformed into an instance of $D$ through polynomial-time reduction.
If any NP-complete problem could be solved with polynomial time complexity, then all problems in NP have polynomial time complexity. Numerous decision forms of optimization issues are considered NP-complete. The class of NP-hard problems is closely associated with NP-complete problems.
@@
Optimization problems on binary input spaces can indeed be transformed into quadratic optimization problems through the following method: Given an integer programming problem with binary decision variables $b_{i}\in\{0,1\}$, $i=1,\dots,n$, this can be rephrased as a quadratic programming problem with continuous decision variables $x_{i}\in\mathbb{R}$ by adding the constraints $(x_{i})(1-x_{i})=1$ for $i=1,\dots,n$. It is clear that continuous optimization issues cannot always be represented as discrete optimization issues. Despite this, some maintain that all problems solved by digital computers are fundamentally discrete and that infinite precision is rarely required in practice. Assuming that operations with infinite precision can be performed in constant time could produce unusual results. For instance, polynomial time algorithms could be formulated for NP-complete problems if the floor function could be computed with infinite precision in polynomial time. Nevertheless, such algorithms are not feasible on a von Neumann architecture with finite precision arithmetic. This scenario underlines the need to consider the computational model alongside complexity outcomes. A noteworthy non-traditional computational model beyond the Turing machine or von Neumann architecture is quantum computing. It is believed that certain algorithms, like prime factorization, can be solved in polynomial time on quantum computers, whereas this is not feasible on classical computers under the assumption that NP is not equal to P.
Finally, in times of growing amounts of decision data, one should not forget that even guarantees of polynomial time complexity can be insufficient in practise. Accordingly, there is a growing interest for problem solvers that require only subquadratic running time. Similar to the construction of the class of $NP$ complete problems, a definition of the class of $3SUM$ complete problems has been constructed by theoretical computer scientists. For this class up to date only slightly better than quadratic running time algorithms are known. A prominent problem from the domain of mathematical programming that belongs to this group is the linear satisfiability problem, i.e. the problem of whether a set of $r$ linear inequality constraints formulated on $n$ continuous variables can be satisfied [34].

### v2

@@
Abstract
Real-world decision and optimization problems, often involve constraints and conflicting criteria. For example, choosing a travel method must balance speed, cost, environmental footprint, and convenience. Similarly, designing an industrial process must consider safety, environmental impact, and cost efficiency. Ideal solutions where all objectives are optimally met are rare; instead, we seek good compromises and aim to avoid lose-lose scenarios. Multicriteria optimization offers computational techniques to compute Pareto optimal solutions, aiding decision analysis and decision making. This reader offers an introduction to this topic and is based on the revised edition of the MSc computer science course lecture ”Multicriteria Optimization and Decision Analysis” at the Leiden Institute of Advanced Computer Science, Leiden University, The Netherlands, conducted from 2007 to 2023. The introduction is organized in a unique didactic manner developed by the authors, starting from more simple concepts such as linear programming and single-point methods, and advancing from these to more difficult concepts such as optimality conditions for nonlinear optimization and set-oriented solution algorithms. Besides, we focus on the mathematical modeling and foundations rather than on specific algorithms, though not excluding the discussion of some representative examples of solution algorithms. Our aim was to make the material accessible to MSc students who do not study mathematics as their core discipline by introducing basic numerical analysis concepts when necessary and providing numerical examples for interesting cases.
@@ Abstract
For several reasons multicriteria optimization and decision making is an exciting field of computer science and operations research. Part of its fascination stems from the fact that in MCO and MCDM different scientific fields are addressed. Firstly, to develop the general foundations and methods of the field, one has to deal with structural sciences, such as algorithmics, relational logic, operations research, and numerical analysis.
@@
  • What are the essential differences between single-objective and multi-objective optimization?
  • How can we classify the solutions? What different types of order relations are used in decision theory, and how are they related to each other?
  • Given a decision model or optimization problem, which formal conditions must be satisfied for the solutions to be optimal?
  • How can we construct algorithms that obtain optimal solutions or approximations to them in an efficient way?
@@
Whenever it comes to decision making in the real world, these decisions will be made by people responsible for it. In order to understand how people come to decisions, and how the psychology of individuals (cognition, individual decision making) and organizations (group decision making) needs to be studied. Questions like the following may arise:
@@
  • What different strategies are used by people to make decisions? How can satisfaction be measured? What strategies are promising in making satisfactory decisions?
@@
  • How do groups of people come to decisions? What are conflicts and how can they be avoided? How do we deal with minority interests in a democratic decision-making process? Can these aspects be integrated into formal decision models?
Moreover, decisions are always related to a real-world problem. Given an application field, we may find very specific answers to the following questions.
@@
  • By which means can we retrieve the values for the criteria (experiments, surveys, function evaluations)? Are there any particular problems with these measurements (dangers and costs) and how to deal with them? What are the uncertainties in these measurements?
@@
  • What are typical decision processes in the field and what implications do they have for the design of decision support systems?
@@ 1.1 Viewing mulicriteria optimization as a task in system design and analysis
The discussion above can be seen as a rough sketch of questions that define the scope of multicriteria optimization and decision-making. However, it needs to be clarified more precisely what is going to be the focus of these lecture notes. For this reason, we want to approach the problem class from the point of view of system design and analysis. Here, with system analysis, we denote the interdisciplinary research field that deals with the modeling, simulation, and synthesis of complex systems.
In addition to experimentation with a physical system, often a system model is used. Today, system models are typically implemented as computer programs that solve (differential) equation systems, simulate interacting automata, or stochastic models. We will also refer to them as simulation models. An example of a simulation model based on differential equations would be the simulation of the fluid flow around an airfoil based on the Navier-Stokes equations. An example of a stochastic system model could be the simulation of a system of elevators, based on some agent-based stochastic model.
@@
In Figure 1.1 different tasks of system analysis based on simulation models are displayed schematically. Modeling means identifying the internal structure of the simulation model. This is done by looking at the relationship between the known inputs and the outputs of the system. In many cases, the internal structure of the system is already known up to a certain granularity and only some parameters need to be identified. In this case we usually speak of calibration of the simulation model instead of modeling. In control theory, the term identification is also common.
Once a simulation-model of a system is given, we can simulate the system, i.e. predict the state of the output variables for different input vectors. Simulation can be used to predict the output for unmeasured input vectors. Usually such model-based predictions are much cheaper than doing the experiment in the real world. Consider, for example, crash test simulations or wind channel simulation. In many cases, such as for future predictions, where time is the input variable, it is even impossible to do the experiments in the physical world. Often the purpose of simulation is also to learn more about the behavior of the systems. In this case systematic experimenting is often used to study effects of different input variables and combinations of them. The field of Design and Analysis of Computer Experiments (DACE) is devoted to such systematic explorations of the behavior of a system.
Finally, we may want to optimize a system: In that case we basically specify what the output of the system should be. We also are given a simulation-model to do experiments with, or even the physical system itself. The relevant open question is how to choose the input variables in order to achieve the desired output. In optimization, we typically want to maximize (or minimize) the value of an output variable.
@@ 1.2 Formal Problem Definitions in Mathematical Programming
The objective function $f$ is a function to be minimized (or maximized [^1] Maximization can be rewritten as minimization by changing the sign of the objective function, that is, replacing $f(\mathbf{x})\rightarrow\max$ with $-f(\mathbf{x})\rightarrow\min$). This is the goal of the optimization. The function $f$ can be evaluated for every point $\mathbf{x}$ in the search space (or decision space). Here the search space is defined by a set of intervals that restrict the range of variables, so-called bounds or box constraints. In addition to this, variables can be integer variables, that is, they are chosen from $\mathbb{Z}$ or subsets of it, or continuous variables (from $\mathbb{R}$). An important special case of integer variables are binary variables which are often used to model binary decisions in mathematical programming.
Whenever inequality and equality constraints are stated explicitly, the search space $\mathcal{X}$ can be partitioned in a feasible search space $\mathcal{X}_{f}\subseteq\mathcal{X}$ and an infeasible subspace $\mathcal{X}-\mathcal{X}_{f}$. In the feasible subspace, all conditions stated in the mathematical programming problem are satisfied. The conditions in the mathematical program are used to avoid constraint violations in the system under design, e.g., the excess of a critical temperature or pressure in a chemical reactor (an example for an inequality constraint) or the keeping of conservation of mass formulated as an equation (an example for an equality constraint). The conditions are called constraints. Due to a convention in the field of operations research, constraints are typically written in a standardized form such that 00 appears on the right-hand side. Equations can easily be transformed into the standard form by means of algebraic equivalence transformations.
Based on this very general definition of problems, we can define several classes of optimization problems by looking at the characteristics of the functions $f$, $g_{i},i=1,\dots,n_{g}$, and $h_{i},i=1,\dots,n_{h}$. Some important classes are listed in Table 1.1.
@@ 1.2.1 Other Problem Classes in Optimization
There are also other types of mathematical programming problem. These are, for instance, based on:
@@
In some cases, the demand that a solution consists of a vector of variables is too restrictive, and instead we can define the search space as some set $\mathcal{X}$. In order to capture also these kind of problems, a more general definition of a general optimization problem can be used:
@@
Another important special case is given, if $\mathcal{X}=\mathbb{R}^{n}$. Such problems are defined as continuous unconstrained optimization problems or simply as unconstrained optimization problems.
For notational convenience, in the following we will refer mainly to the generic definition of an optimization problem given in Equation 1.10, whenever the constraint treatment is not particularly addressed. In such cases, we assume that $\mathcal{X}$ already contains only feasible solutions.
@@ 1.2.2 Multiobjective Optimization
Before proceeding in this direction, it is, however, important to note that many difficulties of solving single objective optimization problems are inherited by the more general class of multiobjective optimization problems. We will therefore first summarize these.
@@ 1.3.1 Problem Difficulty in Continuous Optimization
In continuous optimization, the metaphor of a optimization landscape is often used in order to define problem difficulty. Unlike talking about a function, when using the term (search) landscapes one explicitly requires that the search space be equipped with a neighborhood structure, which could be a metric or a topology. This topology is typically the standard topology in $\mathbb{R}^{n}$ and as a metric typically the Euclidean metric is used.
As we shall discuss with more rigor in Chapter 4, this gives rise to definitions such as local optima, which are points that cannot be improved by replacing them with neighboring points. For many optimum-seeking algorithms, it is difficult to escape from such points or find a good direction (in case of plateaus). If local optima are not also global optima, the algorithm might return suboptimal solutions.
@@
Finally, discontinuities and ruggedness of the landscape make problems difficult to solve for many solvers. Discontinuities are abrupt changes of the function value in some neighborhood. In particular, these cause difficulties for optimization methods that assume the objective function to be continuous, that is, they assume that similar inputs cause similar outputs. A common definition of continuity is that of Lipschitz continuity.
@@
Besides continuity assumptions, also differentiability of the objective function and constraints, convexity, and mild forms of nonlinearity (as given in convex quadratic optimization), as well as limited interaction between variables can make a continuous problem easier to solve. The degree of interaction between variables is given by the number of variables in the term of the objective function: Assume that it is possible to (re)write the optimization problem in the form $\sum_{i=1}^{n}f_{i}(x_{i_{1}},\dots,x_{i_{k(i)}})\rightarrow\max$, then the value of $k(i)$ is the degree of interaction in the $i$-th component of the objective function. In case of continuous objective functions it can be shown that problems with a low degree of interaction can be solved more efficiently in terms of worst-case time complexity[71]. One of the reasons why convex quadratic problems can be solved efficiently is that, given the Hessian matrix, the coordinate system can be transformed by simple rotation in such a way that these problems become decomposable, i.e., $k(i)$ is bounded by $1$.
@@ 1.3.2 Problem Difficulty in Combinatorial Optimization
Having said this, a combinatorial optimization problem can be defined by means of a pseudo-boolean objective function, i.e. $f:\{0,1\}^{n}\rightarrow\mathbb{R}$ and stating the goal $f(\mathbf{x})\rightarrow\min$. Theoretical computer science has developed a rich theory on the complexity of decision problems. A decision problem is the problem of answering a query on input of size $n$ with the answer being either yes or no. In order to relate the difficulty of optimization problems to the difficulty of decision problems, it is beneficial to formulate the so-called decision versions of optimization problems.
@@
A decision problem is said to belong to the class P if there exists an algorithm on a Turing machine222or any in any common programming language operating on infinite memory and not using parallel processing and not assuming constant time complexity for certain infinite precision floating point operations such as the floor function. that solves it with a time complexity that grows at most polynomially with the size $n$ of the input. It belongs to the class NP if a candidate solution $\mathbf{x}$ of size $n$ can be verified (’checked’) with polynomial-time complexity in $n$ (e.g., does it satisfy the formula $f(\mathbf{x})\leq k$ or not). Obviously, the class NP subsumes the class P, but P does not necessarily subsume NP. In fact, the question whether P subsumes NP is the open problem often discussed in theoretical computer science, known as the problem ’P$=$NP?’. Under the assumption ’P $\neq$ NP’, that is that P does not include NP, it is meaningful to introduce the complexity class of NP complete problems:
@@
A decision problem $D$ is NP-complete if it belongs to NP, and every instance of another problem in NP can be transformed into an instance of $D$ through polynomial-time reduction.
If any NP-complete problem could be solved with polynomial-time complexity, then all problems in NP have polynomial time complexity. Numerous decision forms of optimization issues are considered NP-complete. The class of NP-hard problems is closely associated with NP-complete problems.
@@
Optimization problems on binary input spaces can indeed be transformed into quadratic optimization problems through the following method: Given an integer programming problem with binary decision variables $b_{i}\in\{0,1\}$, $i=1,\dots,n$, this can be rephrased as a quadratic programming problem with continuous decision variables $x_{i}\in\mathbb{R}$ by adding the constraints $(x_{i})(1-x_{i})=1$ for $i=1,\dots,n$. It is clear that continuous optimization issues cannot always be represented as discrete optimization issues. Despite this, some maintain that all problems solved by digital computers are fundamentally discrete and that infinite precision is rarely required in practice. Assuming that operations with infinite precision can be performed in constant time could produce unusual results. For instance, polynomial time algorithms could be formulated for NP-complete problems if the floor function could be computed with infinite precision in polynomial time. Nevertheless, such algorithms are not feasible on a von Neumann architecture with finite precision arithmetic. This scenario underlines the need to consider the computational model alongside complexity outcomes. A noteworthy non-traditional computational model beyond the Turing machine or von Neumann architecture is quantum computing. It is believed that certain algorithms, like prime factorization, can be solved in polynomial time on quantum computers, whereas the existence of polynomial-time algorithms is not yet known on classical von Neumann computers or Turing machines [81].
Finally, in times of growing amounts of decision data, one should not forget that even guarantees of polynomial-time complexity can be insufficient in practise. Accordingly, there is a growing interest for problem solvers that require only subquadratic running time. Similarly to the construction of the class of complete NP problems, theoretical computer scientists have constructed a definition of the class of 3SUM-complete problems. For this class up to date only slightly better than quadratic running time algorithms are known, and until very recently it was believed that the quadratic time complexity barrier cannot be surpassed [34]. A prominent problem from the domain of mathematical programming that belongs to this group is the linear satisfiability problem, i.e. the problem of whether a set of $r$ linear inequality constraints formulated on $n$ continuous variables can be satisfied [35].